# Eventually LIL Regret: Almost Sure $\ln\ln T$ Regret for a sub-Gaussian Mixture on Unbounded Data

### v1

@@ 4.1.1 Step 1: proving regret bound along all data sequences
The first two cases in (5) correspond to $t$ such that the unconstrained optimizer $\eta^{*}_{t}$ is in [1,1][-1,1]. To get the bound on $R_{t}$ in (5) in these cases, we show that the mixture wealth $Z_{t}$, which is an integration of a non-negative function over [1,1][-1,1] weighted by a prior $\pi$, is at least the value of the integration over a subset that contains the maximizer $\eta^{*}_{t}$. This subset is chosen large enough so that it contains a significant fraction of the prior mass. Further, it is also small enough so that the value of the function being integrated remains almost constant (and almost equal to the maximum value) in this range.
@@ 4.1.2 Step 2: conditional regret bound on $\mathcal{E}_{\alpha}$
Recall $\mathcal{E}_{\alpha}=\{\sup_{t}\ln Z_{t}\leq\ln(1/\alpha)\}$. Observe that to show (4.1), it suffices to show that the bounds in the middle and last branches of (5) are of the form in (4.1) on $\mathcal{E}_{\alpha}$ (it holds for the first branch with $C=4$).
@@
Rearranging the above inequality, using that ${|S_{t}|}\leq{V_{t}}$, and that on $\mathcal{E}_{\alpha}$, $\sup_{t}\ln Z_{t}\leq\ln\frac{1}{\alpha}$, we get the following on $\mathcal{E}_{\alpha}$ in this case:
@@ 4.1.3 Step 3: stochastic data
Finally, if the data are drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance proxy $(V_{t})_{t\geq 1}$, then $(e^{f_{t}(\eta)})_{t\geq 1}$ is a non-negative super-martingale for every $\eta\in\mathbb{R}$ (Definition 1). Thus, $(Z_{t})_{t\geq 1}$ is also a non-negative super-martingale. The last statement then follows from Ville’s inequality [Ville, 1939]. $\Box$
@@ 5 Conclusions
In this work, we proved deterministic regret bounds for sub-Gaussian mixture processes of the form $Z_{t}$, which mix $\exp\{\eta S_{t}-\eta^{2}V_{t}/2\}$ over $\eta\in\mathbb{R}$. Our regret bounds hold for all sequences of data. However, without any assumption on the data sequence, they can be large (for example, linear in $t$ or $V_{t}$). We then show that for certain sequences of data, specifically along which the mixture $Z_{t}$ remains uniformly bounded, the regret for the mixture using Robbins’ heavy-near-0 prior is $O(\ln\ln V_{t})$ for every $t$. Furthermore, if the data are assumed to be stochastic, then the set of such sequences is large, and hence $O(\ln\ln V_{t})$ holds along many sequences of observations. In fact, we show that regret is $O(\ln\ln V_{t})$ eventually on a set of measure $1$.
@@ Appendix B Proofs from Section 3
$\ln Z_{t}=\frac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})-\frac{\sigma_{0}^{2}S_{t}^{2}}{2(1+\sigma_{0}^{2}V_{t})},$
@@
$\frac{S_{t}^{2}}{1+\sigma_{0}^{2}V_{t}}=\frac{2}{\sigma_{0}^{2}}\left(\ln Z_{t}+\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})\right).$
@@
$R_{t}=\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})+\frac{1}{\sigma_{0}^{2}V_{t}}\left(\ln Z_{t}+\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})\right).$
On $\mathcal{E}_{\alpha}$, since $\ln Z_{t}\leq\ln(1/\alpha)$ for all $t$, we further get
@@
Finally, if $X_{i}$ are iid from $P$, and $(S_{t})_{t\geq 1}$ is sub-Gaussian with variance $(V_{t})_{t\geq 1}$, then for $\eta\in\mathbb{R}$, the process $(e^{f_{t}(\eta)})_{t\geq 1}$ is a non-negative super-martingale (Definition 1). Hence, $(Z_{t})_{t\geq 1}$ is also a non-negative super-martingale. Ville’s inequality [Ville, 1939] then implies that $P[\mathcal{E}_{\alpha}]\geq 1-\alpha$. ∎
@@ Appendix C Proofs from Section 4
$Z_{t}=\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta$ $=e^{L^{\star}_{t}}\int\limits_{-1}^{1}e^{-\frac{V_{t}}{2}(\eta-\eta^{*}_{t})^{2}}\pi(\eta)d\eta\geq e^{L^{\star}_{t}-\frac{1}{2}V_{t}r^{2}}\int\limits_{\begin{subarray}{c}\left\{\eta:|\eta-\eta^{*}_{t}|\leq r\right\}\\ \cap[-1,1]\end{subarray}}\pi(\eta)d\eta.$
@@
$Z_{t}$ $=\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta\geq e^{f_{t}(\eta^{*}_{t})-\frac{V_{t}r^{2}}{2}}\int\limits_{\{|\eta-\eta^{*}_{t}|\leq r\}\cap[-1,1]}\pi(\eta)d\eta.$
@@
$\ln Z_{t}=\ln\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta\geq\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)+\frac{V_{t}}{2}(1-\rho^{2}).$
Now, on $\mathcal{E}_{\alpha}$, since $\sup_{t}\ln Z_{t}\leq\ln(1/\alpha)$, from the above inequality, we must also have
$\frac{V_{t}}{2}(1-\rho^{2})$ $\leq\ln\tfrac{1}{\alpha}-\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)$
$\overset{(a)}{\leq}\ln\tfrac{1}{\alpha}-\ln\tfrac{\rho}{Z_{0}}+\ln\ln\tfrac{c}{1-\rho}+2\ln\ln\ln\tfrac{c}{1-\rho},$
@@
$\ln{~\int\limits_{1-\rho}^{1}\pi(\eta)d\eta~}\geq\ln\tfrac{\rho}{Z_{0}}-\ln\ln\tfrac{c}{1-\rho}-2\ln\ln\ln\tfrac{c}{1-\rho},$
@@
$\ln Z_{t}$ $\geq\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)+\frac{V_{t}}{2}\left((1-2\rho)+2(1-\rho)\alpha_{t}\right)$
@@
On $\mathcal{E}_{\alpha}$, since $\sup_{t}\ln Z_{t}\leq\ln(1/\alpha)$; using this and (C) in the above inequality, and rearranging,
@@ Appendix D The Sub-Gaussian Betting Game
$\pi_{t}(\eta)=\pi(\eta)\frac{\exp(\eta\sum_{i\leq t-1}X_{i}-\eta^{2}(t-1)/2)}{\int_{\mathbb{R}}\exp(\eta\sum_{i\leq t-1}X_{i}-\eta^{2}(t-1)/2)d\eta}.$

### v2

@@ 4.1.1 Step 1: proving regret bound along all data sequences
The first two cases in (7) correspond to $t$ such that the unconstrained optimizer $\eta^{*}_{t}$ is in [1,1][-1,1]. To get the bound on $R_{t}$ in (7) in these cases, we show that the mixture wealth $W_{t}$, which is an integration of a non-negative function over [1,1][-1,1] weighted by a prior $\pi$, is at least the value of the integration over a subset that contains the maximizer $\eta^{*}_{t}$. This subset is chosen large enough so that it contains a significant fraction of the prior mass. Further, it is also small enough so that the value of the function being integrated remains almost constant (and almost equal to the maximum value) in this range.
@@ 4.1.2 Step 2: conditional regret bound on $\mathcal{E}_{\alpha}$
Recall $\mathcal{E}_{\alpha}=\{\sup_{t}\ln W_{t}\leq\ln(1/\alpha)\}$. Observe that to show (4.1), it suffices to show that the bounds in the middle and last branches of (7) are of the form in (4.1) on $\mathcal{E}_{\alpha}$ (it holds for the first branch with $C=4$).
@@
Rearranging the above inequality, using that ${|S_{t}|}\leq{V_{t}}$, and that on $\mathcal{E}_{\alpha}$, $\sup_{t}\ln W_{t}\leq\ln\frac{1}{\alpha}$, we get the following on $\mathcal{E}_{\alpha}$ in this case:
@@ 4.1.3 Step 3: stochastic data
Finally, if the data are drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance proxy $(V_{t})_{t\geq 1}$, then $(e^{f_{t}(\eta)})_{t\geq 1}$ is a non-negative super-martingale for every $\eta\in\mathbb{R}$ (Definition 1). Thus, $(W_{t})_{t\geq 1}$ is also a non-negative super-martingale. The last statement then follows from Ville’s inequality [Ville, 1939]. $\Box$
5 Trade-off between regret and growth rate
In this section, we point out an apparently inherent trade-off between the growth rate of mixture martingales and their rate of accumulating regret: asymptotic optimality in one precludes asymptotic optimality in the other.
When the data is stochastic (i.i.d. from a fixed distribution $Q$), the lower limit $\liminf\nolimits_{t\to\infty}\frac{1}{V_{t}}\ln W_{t}$ is typically $Q$-a.s. a constant, and in that case, that constant is termed as the asymptotic growth rate of $W_{t}$ against $Q$ [Ramdas and Wang, 2025, Chapter 7]. It quantifies the rate at which the wealth grows exponentially under $Q$. In this section, we compare the two wealth processes defined in (3) and (4) at the level of their growth rate and confidence sequence (CS) widths. In contrast to the rest of the paper, this section is purely stochastic. We assume that the data is drawn i.i.d. from a fixed distribution $Q$ that satisfies $\frac{|S_{t}|}{V_{t}}\to m$ as $t\to\infty$, for some $m\in R_{+}$.
For the Gaussian-mixture from Section 3, using the regret bound from (3.1), we have
$\liminf\limits_{t\rightarrow\infty}\frac{\ln W_{t}}{V_{t}}=\liminf\limits_{t\rightarrow\infty}\left(\frac{S^{2}_{t}}{2V^{2}_{t}}-\frac{R_{t}}{V_{t}}\right)\geq\liminf\limits_{t\rightarrow\infty}\frac{S^{2}_{t}}{2V^{2}_{t}}=\frac{m^{2}}{2}.$
However, as we saw in Section 3, this wealth process has a regret that is $O(\ln V_{t})$, and does not achieve $O(\ln\ln V_{t})$ regret.
Let us now consider the mixture wealth process defined in (4) using Robbins’ prior. From Theorem 4.1, we see that
$\limsup\limits_{t\rightarrow\infty}~\frac{R_{t}}{V_{t}}\leq\begin{cases}\limsup\limits_{t\rightarrow\infty}~\frac{1}{2}\left(\frac{|S_{t}|}{V_{t}}-1+\rho\right)^{2},&\text{ if }1<\frac{|S_{t}|}{V_{t}}\\ 0,&\text{ otherwise}.\end{cases}$
Since Robbins’ mixture puts a large prior mass at bets close to 0, it has relatively small mass at larger bets. Therefore, when the data is such that the optimal bet $\frac{|S_{t}|}{V_{t}}$ is far from 0 (in particular >1>1), it lacks sufficient weight to compete with the best-in-hindsight, and hence, suffers large regret in those cases. While the bound above gives a linear upper bound on the regret in these cases, we numerically observe that the regret is indeed linear in such cases. This linear regret, in fact, affects the growth rate of the wealth process, when $m$ is far from 0. To see this, consider $m>1$. In this case, the growth rate for the Robbins’ mixture wealth is smaller than $m^{2}/2$ due to the linear regret suffered in this case. It is $m^{2}/2$ when $m\in[0,1]$.
Next, it is well-known that for a fixed error probability $\alpha\in(0,1)$, the CS derived from the Gaussian mixture wealth (3) are tighter early on (for small $t$) than those obtained using Robbins’ mixture. However, from (4), we can see that they are asymptotically looser with an asymptotic width of $O(\ln V_{t})$, while the latter has $O(\ln\ln V_{t})$ width. The relative tightness of the former (for small $t$) can be explained from the faster growth rate of the corresponding wealth process, as discussed next.
Recall that one way to construct a $(1-\alpha)$-CS for mean is to run, in parallel, a collection of one-sided sequential tests of the null hypothesis that the mean is $m_{0}$, for each $m_{0}\in\mathbb{R}$. At time $t$, the $(1-\alpha)$-CS is the set of all values $m_{0}$ whose corresponding test has not yet stopped (and hence, has not rejected $m_{0}$). For simplicity, consider testing the mean of a Gaussian distribution with known unit variance. To test mean being $m_{0}$, we may use either of the mixture wealth processes in (3) or (4), with $S_{t}=\sum_{i\leq t}(X_{i}-m_{0})$, and $V_{t}=t$. Both wealth processes are non-negative martingales under the null, and therefore yield valid one-sided sequential tests with the stopping time given by
$\tau_{m_{0}}:=\min\left\{t:W_{t}\geq\frac{1}{\alpha}\right\}.$
From above, one can deduce that for fixed $m_{0}$, the stopping time (to reject $m_{0}$) will be smaller if the wealth $W_{t}$ grows faster. Since the growth rate for Robbins’ mixture wealth from (4) is smaller for $m_{0}$ that are “far” from the true mean, the corresponding test rejects $m_{0}$ later than that using Gaussian mixture from (3). However, once the “far-off” $m_{0}$ are excluded, the Robbins’ wealth suffers a smaller regret, and hence has a faster growth rate (in non-asymptotic sense) than the Gaussian mixture. Thus, it rejects the nearby $m_{0}$ faster, leading to narrower CS eventually. In fact, we don’t see the CS derived using Robbins’ mixture wealth process perform poorly in practice. This is because the mean values very far from the true mean are eliminated from both the CSs early on (since we often pick $\alpha=0.05$), since the signal from them is high, but the effect would be more visible for small $\alpha$.
Thus, there appears to be a trade-off between the asymptotic growth rate of the mixture wealth process and the regret incurred. To achieve the smaller $O(\ln\ln V_{t})$ regret when the mean shift is small, the mixture in (4) pays in its asymptotic growth rate against larger mean shifts. To the best of our knowledge, these observations have not been pointed out in the literature.
@@ 6 Conclusions
In this work, we proved deterministic regret bounds for sub-Gaussian mixture processes of the form $W_{t}$, which mix $\exp\{\eta S_{t}-\eta^{2}V_{t}/2\}$ over $\eta\in\mathbb{R}$. Our regret bounds hold for all sequences of data. However, without any assumption on the data sequence, they can be large (for example, linear in $t$ or $V_{t}$). We then show that for certain sequences of data, specifically along which the mixture $W_{t}$ remains uniformly bounded, the regret for the mixture using Robbins’ heavy-near-0 prior is $O(\ln\ln V_{t})$ for every $t$. Furthermore, if the data are assumed to be stochastic, then the set of such sequences is large, and hence $O(\ln\ln V_{t})$ holds along many sequences of observations. In fact, we show that regret is $O(\ln\ln V_{t})$ eventually on a set of measure $1$.
@@
Acknowledgements
SA acknowledges the generous support from the Pratiksha Trust, Bangalore, through the Young Investigator Award, and the DST INSPIRE Faculty Grant IFA24-ENG-389. AR acknowledges funding from NSF grant 2310718 and a Sloan fellowship.
@@ Appendix B Proofs from Section 3
$\ln W_{t}=\frac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})-\frac{\sigma_{0}^{2}S_{t}^{2}}{2(1+\sigma_{0}^{2}V_{t})},$
@@
$\frac{S_{t}^{2}}{1+\sigma_{0}^{2}V_{t}}=\frac{2}{\sigma_{0}^{2}}\left(\ln W_{t}+\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})\right).$
@@
$R_{t}=\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})+\frac{1}{\sigma_{0}^{2}V_{t}}\left(\ln W_{t}+\tfrac{1}{2}\ln(1+\sigma_{0}^{2}V_{t})\right).$
On $\mathcal{E}_{\alpha}$, since $\ln W_{t}\leq\ln(1/\alpha)$ for all $t$, we further get
@@
Finally, if $X_{i}$ are iid from $P$, and $(S_{t})_{t\geq 1}$ is sub-Gaussian with variance $(V_{t})_{t\geq 1}$, then for $\eta\in\mathbb{R}$, the process $(e^{f_{t}(\eta)})_{t\geq 1}$ is a non-negative super-martingale (Definition 1). Hence, $(W_{t})_{t\geq 1}$ is also a non-negative super-martingale. Ville’s inequality [Ville, 1939] then implies that $P[\mathcal{E}_{\alpha}]\geq 1-\alpha$. ∎
@@ Appendix C Proofs from Section 4
$W_{t}=\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta$ $=e^{L^{\star}_{t}}\int\limits_{-1}^{1}e^{-\frac{V_{t}}{2}(\eta-\eta^{*}_{t})^{2}}\pi(\eta)d\eta\geq e^{L^{\star}_{t}-\frac{1}{2}V_{t}r^{2}}\int\limits_{\begin{subarray}{c}\left\{\eta:|\eta-\eta^{*}_{t}|\leq r\right\}\\ \cap[-1,1]\end{subarray}}\pi(\eta)d\eta.$
@@
$W_{t}$ $=\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta\geq e^{f_{t}(\eta^{*}_{t})-\frac{V_{t}r^{2}}{2}}\int\limits_{\{|\eta-\eta^{*}_{t}|\leq r\}\cap[-1,1]}\pi(\eta)d\eta.$
@@
$\ln W_{t}=\ln\int\limits_{-1}^{1}e^{f_{t}(\eta)}\pi(\eta)d\eta\geq\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)+\frac{V_{t}}{2}(1-\rho^{2}).$
Now, on $\mathcal{E}_{\alpha}$, since $\sup_{t}\ln W_{t}\leq\ln(1/\alpha)$, from the above inequality, we must also have
$\frac{V_{t}}{2}(1-\rho^{2})$ $\leq\ln\frac{1}{\alpha}-\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)$
$\overset{(a)}{\leq}\ln\frac{1}{\alpha}-\ln\frac{\rho}{Z_{0}}+\ln\ln\frac{c}{1-\rho}+2\ln\ln\ln\frac{c}{1-\rho},$
@@
$\ln{~\int\limits_{1-\rho}^{1}\pi(\eta)d\eta~}\geq\ln\frac{\rho}{Z_{0}}-\ln\ln\frac{c}{1-\rho}-2\ln\ln\ln\frac{c}{1-\rho},$
@@
$\ln W_{t}$ $\geq\ln\left(\int_{1-\rho}^{1}\pi(\eta)d\eta\right)+\frac{V_{t}}{2}\left((1-2\rho)+2(1-\rho)\alpha_{t}\right)$
@@
On $\mathcal{E}_{\alpha}$, since $\sup_{t}\ln W_{t}\leq\ln(1/\alpha)$; using this and (C) in the above inequality, and rearranging,
@@ Appendix D The Sub-Gaussian Betting Game
$\pi_{t}(\eta)=\frac{\pi(\eta)\exp(\eta\sum_{i\leq t-1}X_{i}-\frac{\eta^{2}}{2}(t-1))}{\int_{\mathbb{R}}\exp(\eta\sum_{i\leq t-1}X_{i}-\frac{\eta^{2}}{2}(t-1))\pi(\eta)d\eta}.$